a student invests two sums of money, $500 and $800, at different interest rates, receiving a total of $55 in interest after one year. the $500 investment receives an interest rate 2% lower than the interest rate for the $800 investment. find the interest rate for each investment.
Answer by mananth(16946)

let $800 investment be at x%
$500 will be @ (x-2)%
..
800*x%+500(x-2)% = 55
multiply by 100
800x+500x-1000=5500
1300x= 5500+1000
1300x= 6500
/1300
x= 5 % for $800 investment
3% for $500 investment